OpenStudy (shesolitt):

PLS HELP I WILL MEDAL AND FAN Given the equation 7(x + 2) = 2x − 1, solve for the variable. Explain each step and justify your process. B. Jeff solved a similar equation below. Is Jeff's solution correct? Explain why or why not. 3x − 2 = 5(x + 2) 3x − 2 = 5x + 10 8x = 8 x = 1

OpenStudy (elitedamosh):

\[7(x+2) = 2x-1\] \[=7x+14=2x-1\] \[7x-2x +14 = 2x-2x -1\] \[5x+14-14=-1-14\] \[5x=-15\] \[x=-3\]

OpenStudy (anonymous):

In other words No the solution is not correct 3x-2=5(x+2) 3x-2=5x+10 3x-5x=10+2 <-- Should be like this -2x=12/-2 x=-6 And: 7(x+2)=2x-1 7x+14=2x-1 7x-2x=-1-14 5x=-15/5 x=-3

OpenStudy (elitedamosh):

B) Jeff's solution is not correct because when he moved(transposed) the 'x' to the other side, he didn't change the sign and when he moved the "2" to the other side he didn't change the sign either

OpenStudy (elitedamosh):

Hope that helped:)

OpenStudy (shesolitt):

THANK YOU SO MUCH YOU GUYS BOTH DESERVE MEDALS!! (: